OpenStudy (anonymous):

solve by factoring n using the priniple of 0producs x^2-x=12

OpenStudy (anonymous):

x^2-x=12 x^2-x -12 =0 (x-4)(x+3)=0 x=4 or -3

OpenStudy (anonymous):

express radicalls in simplest form x^2+3=172

OpenStudy (turingtest):

please post a separate thread for each question